OpenStudy (anonymous):

fraction to lowest terms 25/48 divided by 15/22

OpenStudy (campbell_st):

invert the 2nd fraction and multiply 25/48 x 22/15 = 5/24 x 11/3 = 55/ 72

OpenStudy (anonymous):

is this done for division

OpenStudy (campbell_st):

when dividing a fraction by a fraction, invert the 2nd fraction and multiply

OpenStudy (anonymous):

Oh ok thank you